OpenStudy (anonymous):

7. Graph the following linear system to determine the number of solutions. y= -3x+6 6x+2y=12

OpenStudy (anonymous):

we will get a straight line only

OpenStudy (anonymous):

Same as the last one. Put the second into slope intercept form. 6x+2y=12 2y=-6x+12 y=-3x+6 Now look at your slopes. Again, they are both the same but this time your y intercept is the same too. This means they are the same line. So that means all solutions fit the given set of equations.

OpenStudy (anonymous):

see first eqn y = -3x +6 multiply both sides by 2 2y = -6x + 12 i.e., 6x + 2y = 12